A stone is thrown vertically up with a speed of 14 m/s, and at that moment it had 37 kJ of kinetic energy. What was the mass of the stone?


1
Expert's answer
2021-02-07T19:11:02-0500
"KE=\dfrac{1}{2}mv^2,""m=\dfrac{2KE}{v^2}=\dfrac{2\cdot3.7\cdot10^4\ J}{(14\ \dfrac{m}{s})^2}=377.5\ kg."